OpenStudy (anonymous):

The giant amoeba Chaos chaos is about 10^3 micrometers in diameter. A micrometer is 10^-6 meter. What part of a meter is the diameter? use a positive exponeet to write your awnser

OpenStudy (anonymous):

diameter = 10^3 micrometers. 1 micrometer = 10^(-6) meter. Hence, diameter = 10^3 * 10^(-6) = 10^(3 - 6) = 10^(-3) = 1/(10^3) = 0.001 meter.